# Adaptive NAD: Online and Self-adaptive Unsupervised Network Anomaly Detector

### v3

@@ Abstract
Index Terms:
Anomaly detection, unsupervised learning, interpretability, online learning, dynamic threshold
I Introduction
@@
Primarily, existing unsupervised anomaly detection approaches strongly rely on assumptions of known legitimate [8, 9, 10, 11, 12, 13, 14, 11, 15, 16, 17] and/or even malicious data [18, 19] to support the training of the initial unsupervised model. Moreover, recent existing unsupervised-learning-based research [8, 9, 10, 20, 15] conducts anomaly detection by using a fixed threshold, which tends to fail to detect anomalies when concept drifts exist. To address this, a dynamic threshold is utilized to distinguish normal from anomalies in [11, 21], but suffers from high false positive rates. Additionally, most of existing literature [8, 9, 10, 12, 14, 15, 22, 16] trains their unsupervised models offline without any adaption to drifting patterns, making them not applicable in real-world applications. There is only limited research on online unsupervised anomaly detection, and most of them like [23, 24, 25] consider “online” as the model being trained offline and utilized for real-time inferences without model updating.
@@
Our work. To address the above issues, we propose a general framework, named Adaptive NAD, to interpret and improve online unsupervised deep learning-based anomaly detection in security applications. The high-level goal of Adaptive NAD is to design a novel interpretable online unsupervised framework for anomaly detection that meets several practical security requirements (such as generalization, human-readable, low false positives, and adapting new threats). To this end, we designed two key strategies in Adaptive NAD: (i) interpretable two-layer based unsupervised anomaly detection strategy by leveraging advantages of both supervised and unsupervised models to generate highly confident pseudo labels in highly imbalanced network streaming data. (ii) Online learning strategy for fine-tuning models and regenerating loss distributions aligned with the evolving network events to alleviate the effect of concept drift.
@@
  • $\bullet$
    We compare it with eleven state-of-the-art models for example, GDN [14], DAGMM [22], AOC-IDS [26], DeepFT [11], DeepAID [15], and Online-Offline Framework (OOF) [21] on two classical public anomaly detection datasets (CIC-Darknet2020 [27], CIC-DoHBrw-2020 [28], and Edge-IIoTset [29]) and the results show more than 5.4%, 23.0%, and 3.2% 3335.4%, 23.0%, and 3.2% are the improvements compared with the best state-of-the-art. improvements in SPAUC.
The rest of the paper is outlined as follows: we first give an overview of the system design in Section II, including design goals, system components, and details of the models used in Adaptive NAD. Then, the proposed novel techniques, i.e., interpretable two-layer anomaly detection strategy (ITAD-S), online learning scheme, and threshold calculation technique, are presented in Section III along with a comprehensive complexity and storage analysis of the proposed Adaptive NAD framework. Following it, Section IV describes the experimental setup and results. Existing related literature is discussed in Section VI. Finally, we conclude the paper in Section VII.
II System Design
@@ II-B System Components
  • $\bullet$
    Unsupervised deep Learning Model $M$: This model is used to learn the network traffic using unsupervised deep learning models. In this paper, an unsupervised learning model (i.e., NAD) is utilized to learn the pattern of the normal network traffic (see Section II-C).
  • $\bullet$
@@
  • $\bullet$
    Adaptive Threshold Computation: This model generates two thresholds $T_{1}$ and $T_{2}$ by using Quantiles and Maximum Likelihood Estimation (MLE) (see Section III-C). Thresholds are adaptively changing to incorporate the characteristics of new network traffic to identify normal and abnormal network data.
  • $\bullet$
@@
II-C The Network Anomaly Detection Model (NAD)
@@
(a) NAD Event Model.
(b) RF classifier.
Figure 2: Selected models.
II-D Random Forest
Here, we use the Random Forest algorithm [32] as our supervised machine learning model $M^{\prime}$, which assumes that $X_{1},X_{2},\cdots,X_{M}$ are $M$ attributes. A sample $S$ is denoted as a vector $(x_{1},x_{2},\cdots,x_{M})$, where $x_{i}$ is the observed value of $X_{i}$. We also define the class label $C=\{c_{1},c_{2}\}$ consisting of two different values $c_{1}$ (normal) and $c_{2}$ (abnormal) whose classification depends on the attributes. Hence, the class predicted by the Random Forest method is given as,
@@ II-D Random Forest
III Methodologies
III-A Interpretable Two-layer Anomaly Detection Strategy (ITAD-S)
@@
III-B Online Learning Scheme
@@ 2       $FitDist\leftarrow\text{FitBestDistribution}(LossDB)$;
III-C Adaptive Threshold Calculation Technique
When new network samples arrive, the behavioral patterns of network traffic may change. Consequently, the predefined thresholds for classifying network samples (e.g., normal or abnormal) need to be periodically updated. Therefore, a novel dynamic threshold calculation strategy that automatically learns from the ongoing behaviors and adapts the corresponding thresholds to newer coming data. The initial thresholds $T_{1}$ and $T_{2}$ are calculated based on a limited dataset (e.g., the data from the first day or randomly selected 1% from the dataset). We use these “normal” samples to train the unsupervised learning model. Initially, because we assume all samples are normal, we use only one threshold $T_{1}$. Then, in the second round of training, we use the trained unsupervised model to predict the loss values for new samples. These loss values are divided into normal and abnormal samples using the threshold $T_{1}$. When a sufficient number of abnormal samples are collected (e.g., when the number of abnormal samples reaches 500), we calculate a second threshold $T_{2}$. From the third round of training onward, we follow the process described in Section II for training and updating. In this way, the threshold for the unsupervised model $M$ can be automatically determined for ongoing network anomaly detection.
@@ III-C Adaptive Threshold Calculation Technique
III-D Complexity and storage analysis
@@
IV Evaluation
This section presents the experimental setup and results. The datasets are detailed subsection IV-A. Next, the experiment setup is introduced in subsection IV-B. Finally, in subsection IV-C, we compare the eleven benchmarks and analyze their results.
IV-A Dataset
@@ IV-B Setup
TABLE I: NAD, Fixed-NAD, and Adaptive NAD comparison on CIC-Darknet2020.
@@
TABLE II: NAD, Fixed-NAD, and Adaptive NAD comparison on CIC-DoHBrw-2020.
@@
TABLE III: NAD, Fixed-NAD, and Adaptive NAD comparison on Edge-IIoTset.
@@
TABLE IV: Comparison with classical unsupervised machine learning algorithms on CIC-Darknet2020.
@@
TABLE V: Comparison with classical unsupervised machine learning algorithms on CIC-DoHBrw-2020.
@@
TABLE VI: Comparison with classical unsupervised machine learning algorithms on Edge-IIoTset.
@@
IV-C Results and Analysis
IV-C1 Baseline Comparison Summary
@@
Contribution of each key technique. Table I, Table II, and Table III illustrate the contribution of each key technique, where NAD means that only the unsupervised deep learning $M^{\prime}$ is utilized, and Fixed-NAD is a variant of Adaptive NAD with a constant threshold.
As shown in Table I, Table II, and Table III, Adaptive NAD outperforms other methods. Adaptive NAD achieves detection SPAUC as high as 97.70%, 80.21%, and 99.06% on CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets with about 36.22%, 14.81%, and 43.28% improvement compared to Fixed-NAD and 40.70%, 13.48%, and 45.95% improvement compared with NAD, respectively. In addition, The MDR of NAD is 2.28% lower but its FAR is 15.16% higher than that of Adaptive NAD on the CIC-Darknet2020 dataset, indicating that it balances FAR and MDR less effectively than Adaptive NAD. Then, NAD and Fixed-NAD exhibited a lower FAR on CIC-DoHBrw-2020. However, their MDRs are very high, at 71.55% and 68.63%, respectively. This indicates that they will miss a large number of anomalies. Notably, both the FAR and MDR of the proposed Adaptive NAD are far lower than NAD and Fixed-NAD on the Edge-IIoTset dataset, proving the effectiveness of Adaptive NAD.
IV-C2 Comparison with state-of-the-art
@@
Comparison with unsupervised machine learning models. As shown in Table IV, Table V, and Table VI, Adaptive NAD outperforms classical unsupervised machine learning algorithms on the CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets. It achieves the highest detection SPAUC of 97.70% on the CIC-Darknet2020 dataset with an improvement of approximately 40.69%, 80.21% on the CIC-DoHBrw-2020 dataset with an improvement of approximately 16.33%, and 99.06% on the Edge-IIoTset dataset with an improvement of approximately 24.61%. Adaptive NAD exhibits excellent FAR while maintaining a very low MDR on all the datasets. Specifically, it achieves the lowest FAR of 0.08%, 2.12%, and 0.03% on the CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets, respectively, meaning it generates very few false alarms while accurately detecting anomalies.
Comparison with recent advanced unsupervised algorithms. Table VII, Table VIII, and Table IX shows Adaptive NAD compared with the state-of-the-art unsupervised models in respect of different metrics for the CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets, respectively. Table VII demonstrates Adaptive NAD surpasses all comparison methods across all metrics on the CIC-Darknet2020 dataset with a remarkable SPAUC of over 97.70% and the lowest FAR of 0.08%. This high level of performance highlights Adaptive NAD’s capacity in anomaly detection within dynamic environments. In Table VIII, Adaptive NAD achieves the highest detection SPAUC of 80.21%, improving by approximately 23.04%. It also performs the best in terms of FAR and MDR, exhibiting the lowest MDR and the third-lowest FAR. Although DeepAID and DAGMM have lower FARs, their MDRs are significantly higher, with DeepAID’s MDR as high as 99.76% and DAGMM’s MDR at 83.82%, meaning that these models fail to detect most actual anomalies. Table IX depicts that Adaptive NAD obtains the highest Accuracy, Precision, Recall, F1-score, MDR, and SPAUC on the Edge-IIoTset dataset, indicating its superior anomaly detection capabilities in comprehensive realistic IoT and IIoT cyber security scenarios. Although its FAR is higher than DeepAID, it has a 9.92% lower MDR and 49.65% higher SPAUC compared with DeepAID. Overall, considering all the metrics including accuracy, precision-macro, recall-macro, f1-score-macro, FAR, MDR, and SPAUC, Adaptive NAD is the best design for network anomaly detection.
TABLE VII: Comparison with the state-of-the-art unsupervised models on CIC-Darknet2020.
@@
TABLE VIII: Comparison with the state-of-the-art unsupervised models on CIC-DoHBrw-2020.
@@
TABLE IX: Comparison with the state-of-the-art unsupervised models on Edge-IIoTset.
@@
IV-C3 Comparison with different training settings (e.g., online, initial, and offline)
@@
As shown in Table X, Table XI, and Table XII, online training outperforms other training settings on all three datasets, it achieves SPAUC as high as 97.70%, 80.21%, and 99.06% on CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets with about 11.37%, 16.7%, 47.67% improvement, respectively. In the CIC-Darknet2020 dataset, offline training achieved the lowest MDR. However, it has a relatively high FAR of 27.74%, which means it is prone to generating a significant number of false alarms. Then, in the CIC-DoHBrw-2020 dataset, initial online training and offline training achieved the lowest MDR and FAR, respectively. However, the FAR of initial online training is 7.10%, nearly five percent higher than online training, while the MDR of offline training is as high as 80.10%. Overall, online training balances FAR and MDR better. In the Edge-IIoTset dataset, online training performs best across all evaluation metrics. In particular, it has 47.67% and 26.35% higher SPAUC than initial online training and offline training, proving the proposed Adaptive NAD’s superior performance for network anomaly detection.
@@
TABLE X: Performance (%) of Adaptive NAD on CIC-Darkent2020: comparisons between online training, initial online training, and offline training. $\triangledown$ denotes the performance difference relative to online training. The highest metric performance is bolded.
@@
TABLE XI: Performance (%) of Adaptive NAD on CIC-DoHBrw-2020: comparisons between online training, initial online training, and offline training. $\triangledown$ denotes the performance difference relative to online training. The highest metric performance is bolded.
@@
TABLE XII: Performance (%) of Adaptive NAD on Edge-IIoTset: comparisons between online training, initial online training, and offline training. $\triangledown$ denotes the performance difference relative to online training. The highest metric performance is bolded.

### v4

@@ Abstract
keywords:
Anomaly detection , unsupervised learning , online learning , dynamic threshold
[…]
[1]organization=School of Future Science and Engineering, Soochow University,city=Suzhou, state=Jiangsu, country=China
[…]
[2]organization=the Key Laboratory of Computer Network and Information Integration (Southeast University), Ministry of Education,country=China
[…]
[3]organization=School of Cyber Science and Engineering, Southeast University, city=Nanjing, state=Jiangsu, country=China
11footnotetext: Corresponding author: Jin Wang
1 Introduction
@@
Primarily, existing unsupervised anomaly detection approaches strongly rely on assumptions of known legitimate [8, 9, 10, 11, 12, 13, 14, 11, 15, 16, 17, 18] and/or even malicious data [19, 20] to support the training of the initial unsupervised model. Moreover, existing unsupervised-learning-based research [8, 9, 10, 21, 15] conducts anomaly detection by using a fixed threshold, which tends to fail to detect anomalies when concept drifts exist. To address this, a dynamic threshold is utilized to distinguish normal from anomalies in [11, 22, 23]. Additionally, most of the existing literature [8, 9, 10, 12, 14, 15, 24, 16] trains their unsupervised models offline without any adaptation to drifting patterns, making them not applicable in real-world applications. There is only limited research on online unsupervised anomaly detection, and most of them, like [25, 26, 27], consider “online” as the model being trained offline and utilized for real-time inferences without model updating.
@@
Our work. To address the above issues, we propose a general framework, named Adaptive NAD, to interpret and improve online unsupervised deep learning-based anomaly detection in security applications. The high-level goal of Adaptive NAD is to design a novel interpretable online unsupervised framework for anomaly detection that meets several practical security requirements (such as generalization, low false positives, and adapting to new threats). To this end, we designed two key strategies in Adaptive NAD: (i) an interpretable two-layer based unsupervised anomaly detection strategy by leveraging the advantages of both supervised and unsupervised models to generate highly confident pseudo labels in highly imbalanced network streaming data. (ii) Online learning strategy for fine-tuning models and regenerating loss distributions aligned with the evolving network events to alleviate the effect of concept drift.
@@
  • 3.
    We compare it with eleven state-of-the-art models for example, GDN [14], DAGMM [24], AOC-IDS [28], DeepFT
    [11], DeepAID [15], and Online-Offline Framework (OOF)
    [22] on two classical public anomaly detection datasets (CIC-Darknet2020 [29], CIC-DoHBrw-2020 [30], and Edge-IIoTset [31]) and the results show more than 5.4%, 23.0%, and 3.2% 3335.4%, 23.0%, and 3.2% are the improvements compared with the best state-of-the-art. improvements in SPAUC.
The rest of the paper is outlined as follows: we first give an overview of the system design in Section 2, including design goals, system components, and details of the models used in Adaptive NAD. Then, the proposed novel techniques, i.e., interpretable two-layer anomaly detection strategy (ITAD-S), online learning scheme, and threshold calculation technique, are presented in Section 3 along with a comprehensive complexity and storage analysis of the proposed Adaptive NAD framework. Following it, Section 4 describes the experimental setup and results. Existing related literature is discussed in Section 6. Finally, we conclude the paper in Section 7.
2 System Design
@@ 2.2 System Components
  • 2.
    Unsupervised deep Learning Model $M$: This model is used to learn the network traffic using unsupervised deep learning models. In this paper, an unsupervised learning model (i.e., NAD) is utilized to learn the pattern of the normal network traffic (see Section 2.3).
  • 3.
@@
  • 4.
    Adaptive Threshold Computation: This model generates two thresholds $T_{1}$ and $T_{2}$ by using Quantiles and Maximum Likelihood Estimation (MLE) (see Section 3.3). Thresholds are adaptively changing to incorporate the characteristics of new network traffic to identify normal and abnormal network data.
  • 5.
@@
(a) NAD Event Model.
(b) RF classifier.
Figure 2: Selected models.
2.3 The Network Anomaly Detection Model (NAD)
@@
2.4 Random Forest
Here, we use the Random Forest algorithm [34] as our supervised machine learning model $M^{\prime}$, which assumes that $X_{1},$
$X_{2},\cdots,X_{M}$ are $M$ attributes. A sample $S$ is denoted as a vector $(x_{1},x_{2},\cdots,x_{M})$, where $x_{i}$ is the observed value of $X_{i}$. We also define the class label $C=\{c_{1},c_{2}\}$ consisting of two different values $c_{1}$ (normal) and $c_{2}$ (abnormal) whose classification depends on the attributes. Hence, the class predicted by the Random Forest method is given as,
@@ 2.4 Random Forest
3 Methodologies
3.1 Interpretable Two-layer Anomaly Detection Strategy (ITAD-S)
@@
3.2 Online Learning Scheme
@@ 2       $FitDist\leftarrow\text{FitBestDistribution}(LossDB)$;
3.3 Adaptive Threshold Calculation Technique
When new network samples arrive, the behavioral patterns of network traffic may change. Consequently, the predefined thresholds for classifying network samples (e.g., normal or abnormal) need to be periodically updated. Therefore, a novel dynamic threshold calculation strategy that automatically learns from the ongoing behaviors and adapts the corresponding thresholds to newer coming data. The initial thresholds $T_{1}$ and $T_{2}$ are calculated based on a limited dataset (e.g., the data from the first day or randomly selected 1% from the dataset). We use these “normal” samples to train the unsupervised learning model. Initially, because we assume all samples are normal, we use only one threshold $T_{1}$. Then, in the second round of training, we use the trained unsupervised model to predict the loss values for new samples. These loss values are divided into normal and abnormal samples using the threshold $T_{1}$. When a sufficient number of abnormal samples are collected (e.g., when the number of abnormal samples reaches 500), we calculate a second threshold $T_{2}$. From the third round of training onward, we follow the process described in Section 2 for training and updating. In this way, the threshold for the unsupervised model $M$ can be automatically determined for ongoing network anomaly detection.
@@ 3.3 Adaptive Threshold Calculation Technique
3.4 Complexity and storage analysis
@@
4 Evaluation
This section presents the experimental setup and results. The datasets are detailed subsection 4.1. Next, the experiment setup is introduced in subsection 4.2. Finally, in subsection 4.3, we compare the eleven benchmarks and analyze their results.
4.1 Dataset
@@ 4.2 Setup
Table 1: NAD, Fixed-NAD, and Adaptive NAD comparison on CIC-Darknet2020.
@@
Table 2: NAD, Fixed-NAD, and Adaptive NAD comparison on CIC-DoHBrw-2020.
@@
Table 3: NAD, Fixed-NAD, and Adaptive NAD comparison on Edge-IIoTset.
@@
Table 4: Comparison with classical unsupervised machine learning algorithms on CIC-Darknet2020.
@@
Table 5: Comparison with classical unsupervised machine learning algorithms on CIC-DoHBrw-2020.
@@
Table 6: Comparison with classical unsupervised machine learning algorithms on Edge-IIoTset.
@@
4.3 Results and Analysis
4.3.1 Baseline Comparison Summary
@@
Contribution of each key technique. Table 1, Table 2, and Table 3 illustrate the contribution of each key technique, where NAD means that only the unsupervised deep learning $M^{\prime}$ is utilized, and Fixed-NAD is a variant of Adaptive NAD with a constant threshold.
As shown in Table 1, Table 2, and Table 3, Adaptive NAD outperforms other methods. Adaptive NAD achieves detection SPAUC as high as 97.70%, 80.21%, and 99.06% on CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets with about 36.22%, 14.81%, and 43.28% improvement compared to Fixed-NAD and 40.70%, 13.48%, and 45.95% improvement compared with NAD, respectively. In addition, The MDR of NAD is 2.28% lower but its FAR is 15.16% higher than that of Adaptive NAD on the CIC-Darknet2020 dataset, indicating that it balances FAR and MDR less effectively than Adaptive NAD. Then, NAD and Fixed-NAD exhibited a lower FAR on CIC-DoHBrw-2020. However, their MDRs are very high, at 71.55% and 68.63%, respectively. This indicates that they will miss a large number of anomalies. Notably, both the FAR and MDR of the proposed Adaptive NAD are far lower than NAD and Fixed-NAD on the Edge-IIoTset dataset, proving the effectiveness of Adaptive NAD.
4.3.2 Comparison with state-of-the-art
@@
Comparison with unsupervised machine learning models. As shown in Table 4, Table 5, and Table 6, Adaptive NAD outperforms classical unsupervised machine learning algorithms on the CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets. It achieves the highest detection SPAUC of 97.70% on the CIC-Darknet2020 dataset with an improvement of approximately 40.69%, 80.21% on the CIC-DoHBrw-2020 dataset with an improvement of approximately 16.33%, and 99.06% on the Edge-IIoTset dataset with an improvement of approximately 24.61%. Adaptive NAD exhibits excellent FAR while maintaining a very low MDR on all the datasets. Specifically, it achieves the lowest FAR of 0.08%, 2.12%, and 0.03% on the CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets, respectively, meaning it generates very few false alarms while accurately detecting anomalies.
Comparison with recent advanced unsupervised algorithms. Table 7, Table 8, and Table 9 shows Adaptive NAD compared with the state-of-the-art unsupervised models in respect of different metrics for the CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets, respectively. Table 7 demonstrates Adaptive NAD surpasses all comparison methods across all metrics on the CIC-Darknet2020 dataset with a remarkable SPAUC of over 97.70% and the lowest FAR of 0.08%. This high level of performance highlights Adaptive NAD’s capacity in anomaly detection within dynamic environments. In Table 8, Adaptive NAD achieves the highest detection SPAUC of 80.21%, improving by approximately 23.04%. It also performs the best in terms of FAR and MDR, exhibiting the lowest MDR and the third-lowest FAR. Although DeepAID and DAGMM have lower FARs, their MDRs are significantly higher, with DeepAID’s MDR as high as 99.76% and DAGMM’s MDR at 83.82%, meaning that these models fail to detect most actual anomalies. Table 9 depicts that Adaptive NAD obtains the highest Accuracy, Precision, Recall, F1-score, MDR, and SPAUC on the Edge-IIoTset dataset, indicating its superior anomaly detection capabilities in comprehensive realistic IoT and IIoT cyber security scenarios. Although its FAR is higher than DeepAID, it has a 9.92% lower MDR and 49.65% higher SPAUC compared with DeepAID. Overall, considering all the metrics including accuracy, precision-macro, recall-macro, f1-score-macro, FAR, MDR, and SPAUC, Adaptive NAD is the best design for network anomaly detection.
Table 7: Comparison with the state-of-the-art unsupervised models on CIC-Darknet2020.
@@
Table 8: Comparison with the state-of-the-art unsupervised models on CIC-DoHBrw-2020.
@@
Table 9: Comparison with the state-of-the-art unsupervised models on Edge-IIoTset.
@@
4.3.3 Comparison with different training settings (e.g., online, initial, and offline)
@@
As shown in Table 10, Table 11, and Table 12, online training outperforms other training settings on all three datasets, it achieves SPAUC as high as 97.70%, 80.21%, and 99.06% on CIC-Darknet2020, CIC-DoHBrw-2020, and Edge-IIoTset datasets with about 11.37%, 16.7%, 47.67% improvement, respectively. In the CIC-Darknet2020 dataset, offline training achieved the lowest MDR. However, it has a relatively high FAR of 27.74%, which means it is prone to generating a significant number of false alarms. Then, in the CIC-DoHBrw-2020 dataset, initial online training and offline training achieved the lowest MDR and FAR, respectively. However, the FAR of initial online training is 7.10%, nearly five percent higher than online training, while the MDR of offline training is as high as 80.10%. Overall, online training balances FAR and MDR better. In the Edge-IIoTset dataset, online training performs best across all evaluation metrics. In particular, it has 47.67% and 26.35% higher SPAUC than initial online training and offline training, proving the proposed Adaptive NAD’s superior performance for network anomaly detection.
@@
Table 10: Performance (%) of Adaptive NAD on CIC-Darkent2020: comparisons between online training, initial online training, and offline training. $\triangledown$ denotes the performance difference relative to online training. The highest metric performance is bolded.
@@
Table 11: Performance (%) of Adaptive NAD on CIC-DoHBrw-2020: comparisons between online training, initial online training, and offline training. $\triangledown$ denotes the performance difference relative to online training. The highest metric performance is bolded.
@@
Table 12: Performance (%) of Adaptive NAD on Edge-IIoTset: comparisons between online training, initial online training, and offline training. $\triangledown$ denotes the performance difference relative to online training. The highest metric performance is bolded.